OpenStudy (nali):

find the vertex, focus, and directrix of the parabola and sketch its graph y^2 = -6x

OpenStudy (anonymous):

in standard form, this is a sidways parabola opening to the left... 4px = y^2 where p is the distance from the vertex to the parabola. the vertex is at the origin (0, 0) since 4p=-6, this implies p=-3/2 so your focus has coordinates (-3/2, 0) the directrix is the same distance away but in the opposite direction of the focus. it has equation x=3/2

OpenStudy (nali):

thanku

OpenStudy (anonymous):

yw...:)